OpenStudy (anonymous):

n divided by 12 = 300 n=

geerky42 (geerky42):

Multiply both sides by 12. Try that.

OpenStudy (anonymous):

is it 144?

geerky42 (geerky42):

300*12 = 144?

OpenStudy (anonymous):

i thought it was divisoion

geerky42 (geerky42):

Well, on left side of equation, you are nultiply 12 by n/12. 12s will cancel each other out, leaves n alone.

geerky42 (geerky42):

On other side, you have 300 times 12

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

3600?

geerky42 (geerky42):

So basically you first have \(\dfrac{n}{12}=300\) Multiplying both sides by 12 will get you \(12\dfrac{n}{12}=300\\~\\\cancel{12}\dfrac{n}{\cancel{12}}=300\times12\|~\\n = 300\times12\) Yes, n = 3600

OpenStudy (anonymous):

thank you a bunches :)

geerky42 (geerky42):

Sorry I messed up. but that makes sense?

OpenStudy (anonymous):

yes thnx :)